OpenStudy (anonymous):

Spheres-R-Us ships bags of golf balls. They want each bag to have about 300 balls (plus or minus 6 balls). Choose the absolute value inequality and compound inequality that model the acceptable number of golf balls b in each bag.

OpenStudy (anonymous):

If N is the expected number and you can have x plus or minus, and b is your acceptable range, then N - x <= b <= N + x. Subtracting N from all terms, -x <= b - N <= x, so |b - N| <= x

OpenStudy (anonymous):

what would be the answer though

OpenStudy (anonymous):

About 95% of your work is done here. The question now reverts back to you. How would you substitute and solve? At this forum, answers are not just handed out. We helpers are supposed to show the askers how to set up problems and reason through them. You have an enormous amount of information at hand now. So, again, how would YOU solve it?

OpenStudy (anonymous):

The best I can do at this point, short of handing an answer over, is to show you how absolute value works. Ask a specific question and then we'll go from there.

OpenStudy (anonymous):

As for the answer to this particular problem, I have already identified N and x, "b" remains a variable because each bag of golf balls is potentially different.